# Revealing Nanostructures in High-Entropy Alloys via Machine-Learning Accelerated Scalable Monte Carlo Simulation

### v2

@@ 3 Method
3.1 Theoretical speedup of SMC vs traditional methods
The theoretical speedup of the SMC method as compared to other methods are estimated as follows: For DFT and linear-DFT, we assume that a 100-atoms SCF calculation takes one hour, based on our experience with the MuST-KKR code [66]. The speed of the GNN method (Allegro) is estimated from the $\rm{Li_{3}PO_{4}}$ structure of 421,824 atoms in Ref. [24]. Note that an MC sweep is defined as making an MC trial move over each lattice site. Combining it with the intrinsic $O(N^{3})$ scaling in the DFT method, the total computational cost of DFT then scales as $O(N^{4})$, as signified in Fig. 1 (d). Building on the principles of near-sightedness [67], the LSMS method employs an approximation that confines electron scattering to a local interaction zone. This approximation reduces the computational cost of energy evaluation to linear scaling, therefore the total computational cost for an MC sweep scales approximately as $O(N^{2})$. For a one million-atom system, this reduces the computational cost by a factor of $10^{8}$, as shown in the orange line in Fig. 1 (d). Despite the enhanced scaling behavior, linear-DFT still requires solving the Schrödinger equations explicitly, which contributes a large prefactor to the total computational cost. By replacing the computationally expensive DFT method with machine learning models such as the GNN model, the computational cost can be further reduced by a factor of $10^{8}$, as shown in the green line in Fig. 1 (d). The speedup from the SMC-GPU implementation can be divided into multiple parts: First, the SMC parallelization strategy replaces the energy evaluation of the whole system with calculating the energy changes of the LIZ, which is independent of the system size. This strategy reduces the computational complexity from $O(N^{2})$ to $O(N)$, as illustrated in Fig. 1 (d). Second, the SMC method can simultaneously harness the thousands or tens of thousands of cores in a modern high-performance GPU (e.g. 14,592 FP32 CUDA cores in an NVIDIA H800 GPU). Third, the energy models used in this work are relatively simple compared to the GNN models. Based on the preceding discussion, the total acceleration ratio of the SMC-GPU method relative to a CPU-based DFT method, for a system of $N$ atoms, can be decomposed into contributions from nearsightedness $S_{NS}$, ML acceleration $S_{ML}$, and SMC implementation $S_{SMC-GPU}$, and written as:
@@ 3.2 Implementation of SMC-GPU
    Parallelization implementation: The actual parallelization implementation on GPUs depends on the energy model. For pairwise models such as EPI, the iteration over the $i_{C}$ index is executed in parallel as CUDA threads on the GPU. For more general energy models such as the local_SRO model, the $i_{C}$ index is parallelized in the CUDA blocks dimension, and the calculation of the local energies for each site within LIZ is parallelized with the CUDA threads, as demonstrated in the flowchart in Fig. 7.
@@
    Link-cell size: Using the link-cell algorithm, the one MC sweep (MC trials over every sites) is decomposed into $n_{A}$ sequential steps, with each one of $n_{C}$ parallel moves. For instance, for a $300\times 300\times 300$ bcc supercell (a total of $000300^{3}\times 2=54,000,$ 000 atoms), assuming next-to-nearest-neighbor EPI interaction, then the link-cell length can be chosen as $3\times a$, where a is the lattice constant. The link-cell size $n_{A}$ is $3\times 3\times 3\times 2=54$, and the number of link-cell is $n_{C}=1,000,000$.
@@ 3.3 The EPI and Local_SRO model
where $N$ is the total number of atoms and $\Pi_{m}^{pp^{\prime}}$ is the proportion of $pp^{\prime}$ interaction in the $m$-th neighboring shell. Note that due to the fixed chemical concentration in the canonical system, the single-site term $V_{i}^{p}$ has been absorbed into the constant, and the number of independent EPI parameters is $M(M-1)/2$ for a $M$-component system, which is the reason for the $p^{\prime}<p$ requirement. In practice, the EPI parameters are determined via Bayesian regression [33] or stochastic gradient descent.
@@
is the percentage of $pp^{\prime}$ interactions for the $m$’s coordination shell, and $n_{m}^{pp^{\prime}}$ is the number of pairs with index $(m,p,p^{\prime})$. Note that we limit the quadratic interaction to the same-element pairs to reduce the number of higher-order terms. The third term on the right of Eq. 6 is a $l^{2}$ regularization of the weights of the quadratic term, reflecting our prior that the higher-order interactions should be a small correction to the EPI model. We name it the local_SRO model since $\Pi_{m}^{pp^{\prime}}$ describes the short-range order within the local chemical environment. It should be noted that, unlike EPI, this local SRO model is no longer a linear or even quadratic function of the pairwise interactions. Instead, it can be expressed as a power series (infinite degree polynomial) of the pairwise interactions due to the $(\pi_{m}^{pp^{\prime}})^{2}$ term. This is also the reason we do not refer to it as a “quadratic EPI model”. It is easy to see that the nonlinear interactions can also be modeled with neural networks such as the multilayer perceptrons:
@@ 3.5 MC simulation
FeCoNiAlTi_1M_LocalSRO $72\times 60\times 60\times 4$ $\rm{4\times 4\times 4}$ 0.04703 $10^{5}$ (1000 K)
@@
FeCoNiAlTi_1B_LocalSRO $632\times 632\times 632\times 4$ $\rm{4\times 4\times 4}$ 21.65 $10^{4}$ (1000 K)
@@ 4 Conclusions
  • 1.
    We introduce a Scalable Monte Carlo (SMC) method that overcomes the parallelization bottlenecks inherent in conventional Monte Carlo (MC) simulations. By generalizing the checkerboard algorithm through the introduction of link cells, our approach reduces the computational complexity of an MC sweep from $O(N^{2})$ to $O(N)$. This method is not only applicable to pairwise interactions but also extends to nonlinear local interactions through the introduction of a local interaction zone, making it highly compatible with machine learning (ML)-enhanced atomistic models.
  • 2.
    The GPU-accelerated implementation of the SMC method enables the simulation of atomistic systems exceeding one billion atoms while maintaining the accuracy of density functional theory (DFT). Such unprecedented capability makes it possible for us to directly observe the nanostructures in HEAs, which can be of more than millions of atoms. Understanding the nanoscale evolution of NPs are vital for understanding the origin of the superb mechanical properties in HEAs.
  • 3.
@@
5 Supplementary
5.1 Ising model
Ising model is a fundamental model in statistical mechanics [75]. The study of the Ising model not only helps elucidating a plethora of important physics phenomena, such as ferromagnetic phase transition, renormalization-group, and critical behaviour [5], but also leads to the invention of many widely-used algorithms, such as the Swendsen-Wang algorithm [76], Wang-Landau algorithm [77], and Wolff cluster flipping algorithm [78, 79]. In the simplest form, the Ising Hamiltonian can be written as:
$H(\bm{\sigma})=-\sum_{\langle i,j\rangle}J\sigma_{i}\sigma_{j},$ (7)
where $J$ is the interacting parameter and $\sigma_{i}$ represents the spin at the neighboring lattice site $i$, and $\bm{\sigma}$ denotes a spin configuration in the lattice. Note that the summation is over all sites and their nearest neighbors. For example, in a 2D square lattice, there are four nearest neighbors, and in a 3D square lattice, there are 6 nearest neighbors. It is also straightforward to generalize the above model to include longer-distance pair interactions within a predefined cutoff radius $r_{c}$ in a general lattice. Such Ising-like models can then be applied to study real materials, in which the spin variable $\sigma$ represents the different elements, and the generalized interaction parameters $J_{\sigma_{i}\sigma_{j}}$ can be interpreted as the chemical bonding between atoms of specific elements.
5.2 Statistical mechanics
In statistical mechanics, the probability for the occurrence of configuration $\bm{\sigma}_{k}$ is given by:
$P(\bm{\sigma}_{k})=\frac{e^{-\beta H(\bm{\sigma}_{k})}}{Z},$ (8)
where the partition function is the sum of all the microstate probabilities:
$Z=\sum_{\bm{\sigma}_{k}}e^{-\beta H(\bm{\sigma}_{k})},$ (9)
and $\beta=1/(k_{B}T)$, where $k_{B}$ is the Boltzmann constant, and $T$ is the temperature. After obtaining the partition function, physical observables $\langle\hat{O}\rangle$ such as the expected energy, magnetization, specific heat, and magnetic susceptibility can then be calculated accordingly from
$\langle\hat{O}\rangle=\sum_{k}\frac{O(\sigma_{k})e^{-\beta H(\bm{\sigma}_{k})}}{Z}.$ (10)
5.3 Monte Carlo Simulation Algorithms
A sample following the probability distribution of Eq. 8 can be generated via Markov chain Monte Carlo (MCMC) simulation. For such a purpose, different schemes can be devised, and the Metropolis-Hastings algorithm is among the most commonly used ones. For the Ising model in Eq. 7, the Metropolis-Hasting algorithm starts with proposing flipping the spins of each lattice site. This trial move gives rise to a change of the total energy by $\Delta E$, and the acceptance probability of this trial is given by:
$P=\begin{cases}1,&\Delta E\leq 0\\ \exp(-\beta{\Delta E}),&\Delta E>0.\end{cases}$ (11)
A Monte Carlo step, or sweep, is defined as attempting the above trial exactly one time for each lattice site. A Monte Carlo simulation typically starts with a given number of warm-up steps. These steps are needed for the system to reach thermal equilibrium. After that a measurement of the configuration is made for each step, and the results can be recorded to calculate the physical observables.
The Monte Carlo simulation can be carried out at different conditions, the most common ones are the canonical ensembles, grand-canonical ensembles, and semi-grand canonical ensembles. In a canonical ensemble, the number of particles and the temperature are kept fixed. In the grand-canonical ensemble, the chemical potential is kept fixed, and the number of particles is determined by the equilibrium condition. The semi-grand canonical is a technique to fix the concentration of different elements by adjusting the value of the chemical potentials of each element. In this work, we focus on the canonical ensemble, which is relatively straightforward to implement. In a canonical Monte Carlo simulation, the first thing to note is that the flipping trial in the Ising model is replaced with a swapping trial, in which a random pair of sites is chosen randomly to switch the particles.
5.4 Parallelization opportunities in Monte Carlo simulation
Compared to molecular dynamics, Monte Carlo simulation is generally considered to be difficult for large-scale parallelization due to the sequential nature of the Metropolis updating. Nevertheless, there are still some parallelization opportunities that can be exploited. These opportunities can be generally divided into three groups:
  • 1.
    Temperature parallelism: the simulations at different temperatures are independent, therefore they are “embarrassingly parallel”. A closely related case is the replica exchange method, in which the configurations from different temperatures are constantly exchanged in the simulation process, with a policy satisfying detailed balance condition, in order to enhance the speed for reaching thermal equilibrium.
  • 2.
    Move parallelism: for two sites far-away from each other, the flipping or swapping move at one site won’t affect the energy of the other, due to the short-range nature of the interactions. This property can therefore presents a parallelization opportunity. Similarly, it is also possible to split a huge lattice to smaller ones, as long as the “ghost cells” at the boundary are taken care of.
  • 3.
    Sub-move parallelism: For calculating the energy change due to a single MC move, some fine-grain parallelization opportunities exist, depending on the form of the effective Hamiltonian. This is particularly true for the models based on neural networks, where a large number of matrix-multiplication operations present excellent opportunities for acceleration with highly optimized math libraries.
In this work, we mainly focus on move-parallelism, which is one of the bottleneck for large-scale MC simulations. For sub-move parallelism, we make use of the fact that, in the general case, all the local energies of sites within a LIZ need to be evaluated, and the calculations of these local energies are independent of each other.
@@ 5.4 Parallelization opportunities in Monte Carlo simulation
The data that support the findings in this study will be published upon the acceptance of this manuscript.
@@ Acknowledgements
The work of Xianglin Liu and Fanli Zhou was supported by the National Natural Science Foundation of China under Grant 12404283. The work of Yonghong Tian was supported by the National Natural Science Foundation of China under Grant Grant 62425101, 62088102. This work was also supported by the Pengcheng Laboratory Key Project (PCL2021A13) and utilized the computing resources of Pengcheng Cloud Brain.
@@
Xianglin Liu: Conceptualization, Methodology, Software, Funding Acquisition, Project Administration, Data Curation, Formal Analysis, Investigation, Visualization, Writing - original draft. Kai Yang: Software, Data Curation, Formal Analysis, Investigation. Fanli Zhou: Visualization, Writing - review & editing. Yongxiang Liu: Writing –review & editing. Dengdong Fan: Writing –review & editing. Pengxiang Xu: Resources, Funding Acquisition, Project Administration, Writing –review & editing. Zongrui Pei Formal Analysis, Writing -review & editing Yonghong Tian: Funding Acquisition, Writing –review & editing.

### v3

@@ 3 Method
3.1 Theoretical speedup of SMC-X vstraditional methods
The theoretical speedup of the SMC-X method as compared to other methods are estimated as follows: For DFT and linear-DFT, we assume that a 100-atoms SCF calculation takes one hour, based on our experience with the MuST-KKR code [80]. The speed of the GNN method (Allegro) is estimated from the $\rm{Li_{3}PO_{4}}$ structure of 421,824 atoms in Ref. [24]. Note that an MC sweep is defined as making an MC trial move over each lattice site. Combining it with the intrinsic $O(N^{3})$ scaling in the DFT method, the total computational cost of DFT then scales as $O(N^{4})$, as signified in Fig. 1 (d). Building on the principles of near-sightedness [81], the LSMS method employs an approximation that confines electron scattering to a local interaction zone. This approximation reduces the computational cost of energy evaluation to linear scaling, therefore the total computational cost for an MC sweep scales approximately as $O(N^{2})$. For a one million-atom system, this reduces the computational cost by a factor of $10^{8}$, as shown in the orange line in Fig. 1 (d). Despite the enhanced scaling behavior, linear-DFT still requires solving the Schrödinger equations explicitly, which contributes a large prefactor to the total computational cost. By replacing the computationally expensive DFT method with machine learning models such as the GNN model, the computational cost can be further reduced by a factor of $10^{8}$, as shown in the green line in Fig. 1 (d). The speedup from the SMC-GPU implementation can be divided into multiple parts: First, the SMC-X parallelization strategy replaces the energy evaluation of the whole system with calculating the energy changes of the LIZ, which is independent of the system size. This strategy reduces the computational complexity from $O(N^{2})$ to $O(N)$, as illustrated in Fig. 1 (d). Second, the SMC-X method can simultaneously harness the thousands or tens of thousands of cores in a modern high-performance GPU (e.g. 14,592 FP32 CUDA cores in an NVIDIA H800 GPU). Third, the energy models used in this work are relatively simple compared to the GNN models. Based on the preceding discussion, the total acceleration ratio of the SMC-GPU method relative to a CPU-based DFT method, for a system of $N$ atoms, can be decomposed into contributions from nearsightedness $S_{NS}$, ML acceleration $S_{ML}$, and SMC-X implementation $S_{SMC-GPU}$, and written as:
@@ 3.2 Implementation of SMC-GPU
    Parallelization implementation: The actual parallelization implementation on GPUs depends on the energy model. For pairwise models such as EPI, the iteration over the $i_{C}$ index is executed in parallel as CUDA threads on the GPU. For more general energy models such as the qSRO model, the $i_{C}$ index is parallelized in the CUDA blocks dimension, and the calculation of the local energies for each site within LIZ is parallelized with the CUDA threads, as demonstrated in the flowchart in Fig. 7.
@@
    Link-cell size: Using the link-cell algorithm, the one MC sweep (MC trials over every sites) is decomposed into $n_{A}$ sequential steps, with each one of $n_{C}$ parallel moves. For instance, for a $300\times 300\times 300$ bcc supercell (a total of $300^{3}\times 2=54,000,000$ atoms), assuming next-to-nearest-neighbor EPI interaction, then the link-cell length can be chosen as $3\times a$, where a is the lattice constant. The link-cell size $n_{A}$ is $3\times 3\times 3\times 2=54$, and the number of link-cell is $n_{C}=1,000,000$.
@@ 3.3 The EPI and qSRO model
where $N$ is the total number of atoms and $\Pi_{m}^{pp^{\prime}}$ is the proportion of $pp^{\prime}$ interaction in the $m$-th neighboring shell. Note that due to the fixed chemical concentration in the canonical system, the single-site term $V_{i}^{p}$ has been absorbed into the constant, and the number of independent EPI parameters is M(M1)/2M(M-1)/2 for a $M$-component system, which is the reason for the $p^{\prime}<p$ requirement. In practice, the EPI parameters are determined via Bayesian regression [33] or stochastic gradient descent.
@@
is the percentage of $pp^{\prime}$ interactions for the $m$’s coordination shell, and $n_{m}^{pp^{\prime}}$ is the number of pairs with index $(m,p,p^{\prime})$. Note that we limit the quadratic interaction to the same-element pairs to reduce the number of higher-order terms. The third term on the right of Eq. 6 is a $l^{2}$ regularization of the weights of the quadratic term, reflecting our prior that the higher-order interactions should be a small correction to the EPI model. We name it the qSRO model since $\Pi_{m}^{pp^{\prime}}$ describes the short-range order within the local chemical environment. It should be noted that, unlike EPI, this qSRO model is no longer a linear or even quadratic function of the pairwise interactions. Instead, it can be expressed as a power series (infinite degree polynomial) of the pairwise interactions due to the $(\pi_{m}^{pp^{\prime}})^{2}$ term. It is easy to see that the nonlinear interactions can also be modeled with neural networks such as the multilayer perceptrons:
@@ 3.5 MC simulation
FeCoNiAlTi_1M_qSRO $72\times 60\times 60\times 4$ $\rm{4\times 4\times 4}$ 0.04703 $10^{5}$ (1000 K)
@@
FeCoNiAlTi_1B_qSRO $632\times 632\times 632\times 4$ $\rm{4\times 4\times 4}$ 21.65 $10^{4}$ (1000 K)
@@ 4 Conclusions
We introduce a Scalable Monte Carlo (SMC-X) method that overcomes the parallelization bottlenecks inherent in conventional Monte Carlo (MC) simulations. By generalizing the checkerboard algorithm through the introduction of link cells, our approach reduces the computational complexity of an MC sweep from $O(N^{2})$ to $O(N)$. This method is not only applicable to pairwise interactions but also extends to nonlinear local interactions through the introduction of a local interaction zone, making it highly compatible with machine learning (ML)-enhanced atomistic models.
The GPU-accelerated implementation of the SMC-X method enables the simulation of atomistic systems exceeding one billion atoms while maintaining the accuracy of density functional theory (DFT). Such unprecedented capability makes it possible for us to directly observe the nanostructures in HEAs, which can be of more than millions of atoms. Understanding the nanoscale evolution of NPs are vital for understanding the origin of the superb mechanical properties in HEAs.
@@
Code availability statement
Executable binaries or a containerized version of the code are available from the authors upon reasonable request.
@@
The data supporting the findings of this study, including the simulated configurations in .xyz format, Python scripts for data analysis, andC++ code for nanoparticle extraction, are available at: https://github.com/xianglil/SMCX-Supplementary
@@ Acknowledgements
The work of X. Liu and F. Zhou was supported by the National Natural Science Foundation of China under Grant 12404283. The work of Y. Tian was supported by the National Natural Science Foundation of China under Grant Grant 62425101, 62088102. This work was also supported by the Pengcheng Laboratory Key Project (PCL2021A13) to utilized the computing resources of Pengcheng Cloud Brain.
X. Liu gratefully acknowledges Professor Robert Swendsen for his guidance and foundational training in the fascinating field of Monte Carlo simulation during his graduate studies at Carnegie Mellon University.
@@
X.L. proposed the SMC method, designed the workflow, calculated the DFT data, trained the energy model, interpreted the data, acquiredthe funding, supervised the work, and wrote the main manuscript text. K.Y. wrote the SMC-GPU code based on X.L.’s prototype, and ranthe MC simulations. F.Z. contributed to Fig. 1d , 4a, 4b and Table 1. Z.P. co-analyzed the data. P.X. co-analyzed the data, acquired the computing resources, and co-supervised the work. Y.T. acquired the funding. All authors reviewed the manuscript.